Solve the equation by completing the square
4x^2 + 16x + 7 = 0

Answer by stanbon(75887)

Solve the equation by completing the square
4x^2 + 16x + 7 = 0
4(x^2+4x+? = -7 + 4*?
4(x^2+4x+4)= -7 + 4*4
4(x+2)^2 = 9
(x+2)^2 = 9/4
x+2 = 9/2 or x+2 = -9/2
x = -2+(9/2) or x = -2 -(9/2)
x = 5/2 or x = -13/2
========================
Cheers,
Stan H.